# Vacuum bubble and fissure formation in collective motion with competing attractive and repulsive forces Olivia Clifton1, Angel Chavez2, Antonio Madrigal3, Annie Warren2, Paige Yeung4, and Arnd Scheel2

### v1

@@ 2.3 Vacuum Formation: Hexagonal Lattice Symmetry
$\langle\mathbf{e}_{i},\mathbf{e}_{j}^{\star}\rangle=\left\{\begin{array}[]{ll}0,&i\neq j\\ 1,&i=j\end{array}\right.,\qquad\qquad\langle\mathbf{e}_{i},\mathbf{e}_{j}\rangle=\left\{\begin{array}[]{ll}0,&i\neq j\\ 1,&i=j\end{array}\right.$. (2.19)
@@
$B_{1}$$B_{2}$$B_{1}$$B_{2}$$B_{1}$$B_{2}$$\pi$$\frac{2\pi}{\sqrt{3}}$$\mathbf{e}_{2}^{\star}$$\mathbf{e}_{1}^{\star}$$\mathbf{e}_{1}$$\mathbf{e}_{2}$$\mathbf{e}_{1}^{\star}-\mathbf{e}_{2}^{\star}$xxyy00
@@
Figure 2.2: Inner figures (blue and green): An illustration of the vacuum bubbles on the hexagonal lattice. Left, triangles; right, hexagons. Blue regions represent areas of vacuum. In the case of triangles, bubbles are initially approximately circular, as pictured here; the triangular corrections become more apparent as $\widetilde{\mu}$} increases. All bubbles have (inner) radius $\ell$. Outer figures (pink and blue): example boundaries of triangle and hexagon vacuum regions from numerics, for various $\widetilde{\mu}$ (size not to scale). Triangles: $\widetilde{\mu}$= 1e-5, 5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2e-3. Hexagons: $\widetilde{\mu}$= 5e-5, .001, .005, .01, .03, .07.
@@
Let $V(x,y)$ as in (2.20), and ${\mu}=\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}$, with $\widetilde{\mu}$} sufficiently small. Then solutions $u$ to (2.4)-(2.6) with vacuum are of the form
$u(x,y)=\left(A_{0}+\frac{1}{3}\left(A_{1}\cos\left(\mathbf{e}_{1}^{\star}\cdot(x,y)\right)+B_{1}\cos\left(\mathbf{e}_{2}^{\star}\cdot(x,y)\right)+C_{1}\cos\left((\mathbf{e}_{1}^{\star}-\mathbf{e}_{2}^{\star})\cdot(x,y)\right)\right)\right)_{+},$ (2.21)
@@
    The half-width of the fissures $\ell=$ and area $\mathcal{A}$ of vacuum region are
    $\ell=\frac{3^{5/6}}{2^{4/3}}\pi{\widetilde{\mu}}^{1/3}+\mathcal{O}({\widetilde{\mu}}^{\frac{2}{3}}),\qquad\mathcal{A}=\frac{3^{5/6}}{2^{1/3}}\pi^{2}{\widetilde{\mu}}^{1/3}+\mathcal{O}({\widetilde{\mu}}^{\frac{2}{3}})$
@@
    $A_{0}=1-\frac{2\sqrt{3}\pi^{2}}{3}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{\frac{3}{2}}),\qquad A_{1}=\frac{2}{3}+\frac{4\pi^{\frac{3}{2}}}{3\sqrt{3}}{\widetilde{\mu}}^{\frac{1}{2}}-\frac{4\pi^{2}}{3\sqrt{3}}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{\frac{3}{2}}).$
@@
    $A_{0}=1-\frac{\pi^{2}}{18\sqrt{3}}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{\frac{3}{2}}),\qquad A_{1}=\frac{1}{3}+\frac{\sqrt{2\pi^{3}}}{9}{\widetilde{\mu}}^{\frac{1}{2}}-\frac{\pi^{2}}{18\sqrt{3}}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{\frac{3}{2}}).$
@@
(hexagons) $A_{\textrm{hex}}=\sqrt{6}\pi^{\frac{5}{2}}{\widetilde{\mu}}^{\frac{1}{2}},$
(triangles) $A_{\textrm{tri}}=4\sqrt{3}\pi^{\frac{5}{2}}{\widetilde{\mu}}^{\frac{1}{2}}=2\sqrt{2}A_{\textrm{hex}}$
The triangle solutions thus possess not only a greater maximum density than the hexagons, but a greater area of vacuum.
We shall prove existence and establish expansions for triangles and hexagons in the next two sections. Existence and expansion for fissures again follows from the one-dimensional case [13]. Similar to the square case, Remark 2.1, the branches found here have maximal isotropy. There do not appear to be solution branches with submaximal isotropy.
@@
One can easily plot the sublevel sets of the potential $V$ and its negative to identify the shape of vacuum regions; compare Fig. 1.1. In the case of square lattices, the vacuum bubbles are asymptotically discs with small square-symmetric corrections. This is also apparent from the formula (2.11). In the case of the hexagonal lattice, the corrections to the shape are triangular and hexagonal, respectively; see again Fig. 1.1 and 2.2).
@@ 2.3.1 Proof of Theorem 2 (ii) — triangles
where the minus sign is taken for bubbles labeled $B_{1}$ and the plus sign is used for bubbles labeled $B_{2}$ (see Fig. 2.2). We identify the bubble radius $\ell$ as the solution to (2.23), and further scale $\ell=\ell_{1}z_{1}$. In this scaling, after some algebraic manipulation, (2.22) becomes
$0=-\frac{3}{2(1+a_{0}+z_{1}^{2})}+\frac{\ell_{1}^{2}}{2}-\frac{z_{1}^{2}\ell_{1}^{4}}{24}+\mathcal{O}(r^{5})+Q(\theta),$ (2.23)
@@
$\ell_{1}(z_{1},a_{0},\theta)$ $=\sqrt{3}-\frac{1}{\sqrt{3}}z_{1}^{2}-\frac{\sqrt{3}}{2}a_{0}+\mathcal{O}\left(z_{1}^{4}+a_{0}^{2}\right).$
Again, we can use this expression for $\ell(z_{1},a_{0},\theta)$ to solve for $a_{0}$ using the average-density condition (2.5). Substituting, we obtain
$0$ $=2\sqrt{3}\pi^{2}\cdot a_{0}-\iint_{\Omega_{0}^{c}}\left(\left(1+a_{0}\right)+\frac{2}{3}(1+a_{0}+z_{1}^{2})\left(2\cos(x)\cos\left(\frac{y}{\sqrt{3}}\right)+\cos\left(\frac{2y}{\sqrt{3}}\right)\right)\right)dydx$
$=2\sqrt{3}\pi^{2}\cdot a_{0}-2\int_{0}^{2\pi}\int_{0}^{z_{1}\ell_{1}(z_{1},a_{0},\theta)}\left[(1+a_{0})\frac{r^{2}}{3}+z_{1}^{2}\left(-1+\frac{r^{2}}{3}\right)+\mathcal{O}(r^{4})\right]r\ drd\theta$
$=2\sqrt{3}\pi^{2}\cdot a_{0}+3\pi(1-a_{0})z_{1}^{4}+\mathcal{O}\left(z_{1}^{6}\right),$
@@
Lastly, we use (2.4) to relate these quantities to $\widetilde{\mu}$}, using polar coordinates for the integral over the bubbles:
$0=$ $V(x,y)\left(A_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(A_{1}\sqrt{3}\pi^{2})\right)$
$+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}V(x-\xi,y-\eta)\left(A_{0}+A_{1}\cdot V(\xi,\eta)\right)d\xi d\eta$
$=$ $-\sqrt{3}{\widetilde{\mu}}\pi^{2}A_{1}+2\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\int_{0}^{2\pi}\int_{0}^{z_{1}\ell_{1}}\left(\frac{1}{2}z_{1}^{2}-\frac{1}{6}r^{2}+\mathcal{O}(r^{4},r^{2}z_{1}^{2})\right)r\ drd\theta$
@@
As in the square symmetry case for bubbles, we find that $z_{1}$ scales like ${\widetilde{\mu}}^{\frac{1}{4}}$, and we obtain a unique $z_{1}=z_{1}({\widetilde{\mu}}^{\frac{1}{4}})$, for ${\widetilde{\mu}}^{\frac{1}{4}}$ sufficiently small, by the implicit function theorem, with
$z_{1}=\left(\frac{4\pi^{3}}{3}{\widetilde{\mu}}\right)^{\frac{1}{4}}+\mathcal{O}({\widetilde{\mu}}^{\frac{1}{2}}).$ (2.25)
@@ 2.3.2 Proof of Theorem 2 (iii) — hexagons
We now turn to the case $A_{1}>0$. and proceed analogously. We scale
$A_{0}$ $=1+a_{0},\qquad A_{1}=\frac{1}{3}\left(1+a_{0}+z_{1}^{2}\right),\qquad a_{0}=a_{0}(z_{1})$
for a small parameter $z_{1}$. Setting the solution equal to 0 in the free-boundary equation (2.6), expanding in polar coordinates, yields
@@
$0$ $=-2\sqrt{3}\pi^{2}\cdot a_{0}+\iint_{\Omega_{0}^{c}}\left(\left(1+a_{0}\right)-\frac{1}{3}(1+a_{0}+z_{1}^{2})\left(2\cos(x)\cos\left(\frac{y}{\sqrt{3}}\right)+\cos\left(\frac{2y}{\sqrt{3}}\right)\right)\right)dydx$
@@
$0=$ $V(x,y)\left(-A_{1}+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(A_{1}\sqrt{3}\pi^{2})\right)$
$+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}V(x-\xi,y-\eta)\left(A_{0}-A_{1}V(\xi,\eta)\right)dxdy$
@@
$=$ $\frac{{\widetilde{\mu}}\pi^{2}}{\sqrt{3}}\left(1+z_{1}^{2}-\frac{\sqrt{3}}{4\pi}z_{1}^{4}\right)+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\left(-\frac{3\pi}{2}z_{1}^{4}+\mathcal{O}\left(z_{1}^{6}\right)\right).$
Again, we find that $z_{1}$ scales like ${\widetilde{\mu}}^{\frac{1}{4}}$, and we obtain a unique $z_{1}=z_{1}({\widetilde{\mu}}^{\frac{1}{4}})$, for ${\widetilde{\mu}}^{\frac{1}{4}}$ sufficiently small, by the implicit function theorem, with
$z_{1}=\left(\frac{2\pi^{3}}{3}{\widetilde{\mu}}\right)^{\frac{1}{4}}+\mathcal{O}({\widetilde{\mu}}^{\frac{1}{2}}).$ (2.30)

### v2

@@ 2.3 Vacuum Formation: Hexagonal Lattice Symmetry
$\langle\mathbf{e}_{i},\mathbf{e}_{j}^{\star}\rangle=\left\{\begin{array}[]{ll}0,&i\neq j\\ 1,&i=j\end{array}\right$., (2.17)
@@
$G_{1}$$G_{2}$$G_{1}$$G_{2}$$G_{1}$$G_{2}$$\pi$$\frac{2\pi}{\sqrt{3}}$$\mathbf{e}_{2}^{\star}$$\mathbf{e}_{1}^{\star}$$\mathbf{e}_{1}$$\mathbf{e}_{2}$$\mathbf{e}_{1}^{\star}-\mathbf{e}_{2}^{\star}$xxyy00
@@
Figure 2.2: Inner figures (blue and green): An illustration of the vacuum bubbles on the hexagonal lattice. Left, triangles; right, hexagons. Blue regions represent areas of vacuum. In the case of triangles, bubbles are initially approximately circular, as pictured here; the triangular corrections become more apparent as $\widetilde{\mu}$} increases. All bubbles have (inner) radius $\ell$. Outer figures (pink and blue): example boundaries of triangle and hexagon vacuum regions from numerics, for various $\widetilde{\mu}$ (size not to scale with inner diagram). Triangles: $\widetilde{\mu}$= 1e-5, 5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2e-3. Hexagons: $\widetilde{\mu}$= 5e-5, .001, .005, .01, .03, .07.
@@
Let $V(x,y)$ as in (2.18), and ${\mu}=\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}$, with $\widetilde{\mu}$} sufficiently small. Then solutions $u$ to (2.4)-(2.6), invariant under point reflections u(x,y)=u(x,y)u(x,y)=u(-x,-y) and with vacuum regions, are of the form
$u(x,y)=\left(A_{0}+A_{1}\cos\left(\mathbf{e}_{1}^{\star}\cdot(x,y)\right)+B_{1}\cos\left(\mathbf{e}_{2}^{\star}\cdot(x,y)\right)+C_{1}\cos\left((\mathbf{e}_{1}^{\star}-\mathbf{e}_{2}^{\star})\cdot(x,y)\right)\right)_{+},$ (2.19)
@@
    The half-width of the fissures $\ell$ and area $\mathcal{A}$ of vacuum region are
    $\ell=\frac{3^{5/6}}{2^{4/3}}\pi{\widetilde{\mu}}^{1/3}+\mathcal{O}({\widetilde{\mu}}^{2/3}),\qquad\mathcal{A}=\frac{3^{5/6}}{2^{1/3}}\pi^{2}{\widetilde{\mu}}^{1/3}+\mathcal{O}({\widetilde{\mu}}^{2/3}).$
@@
    $A_{0}=1-\frac{2\pi^{2}}{\sqrt{3}}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{{3}/{2}}),\qquad A_{1}=\frac{2}{3}+\frac{4\pi^{3/2}}{3\sqrt{3}}{\widetilde{\mu}}^{{1}/{2}}-\frac{4\pi^{2}}{3\sqrt{3}}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{{3}/{2}}).$
@@
    $A_{0}=1-\frac{\pi^{2}}{18\sqrt{3}}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{{3}/{2}}),\qquad A_{1}=-\frac{1}{3}-\frac{\sqrt{2\pi^{3}}}{9}{\widetilde{\mu}}^{1/2}+\frac{\pi^{2}}{18\sqrt{3}}{\widetilde{\mu}}+\mathcal{O}({\widetilde{\mu}}^{{3}/{2}}).$
@@
(hexagons) $A_{\textrm{hex}}=\sqrt{6}\pi^{{5}/{2}}{\widetilde{\mu}}^{1/2},$
(triangles) $A_{\textrm{tri}}=4\sqrt{3}\pi^{{5}/{2}}{\widetilde{\mu}}^{1/2}=2\sqrt{2}A_{\textrm{hex}}.$
The triangle solutions thus possess not only a larger maximum density than the hexagons, but a larger area of vacuum.
We shall prove existence and establish expansions for triangles and hexagons in the next two sections. Existence and expansion for fissures again follow from the one-dimensional case [15]. Similar to the square case, Remark 2.1, the branches found here have maximal isotropy.
@@
One can easily plot the sublevel sets of the potential $V$ and its negative to identify the shape of vacuum regions; compare Fig. 1.1. In the case of square lattices, the vacuum bubbles are asymptotically discs with small square-symmetric corrections. This is also apparent from the formula (2.11). In the case of the hexagonal lattice, the corrections to the shape are triangular and hexagonal, respectively; see again Fig. 1.1 and 2.2.
Remark 2.4 (Maximal Isotropy — hexagonal lattice).
Similar to the case of the square lattice, we also investigated the possibility of solutions with submaximal isotropy. We did not find such solutions either in the finite-rank approximation or numerically including small viscosity. We particularly investigated the possibility of triangles, which notably do not possess the point reflection symmetry $(x,y)\mapsto(-x,-y)$, with isotropy $D_{3}$, submaximal and did not find such solutions in this finite-rank approximation. We also did not find those numerically, nor does there appear to exist a secondary branch of triangles as sometimes observed in bifurcations with this symmetry group of the hexagonal lattice; see Fig. 4.3
@@ 2.3.1 Proof of Theorem 2 (ii) — triangles
where the minus sign is taken for bubbles labeled $G_{1}$ and the plus sign is used for bubbles labeled $G_{2}$ (see Fig. 2.2). We identify the bubble radius $\ell$ as the solution $r$ to (2.20), and further scale $\ell=\ell_{1}z_{1}$. In this scaling, after some algebraic manipulation, (2.20) becomes
$0=-\frac{3}{2(1+a_{0}+z_{1}^{2})}+\frac{\ell_{1}^{2}}{2}+z_{1}\ell_{1}^{3}Q(\theta)-\frac{z_{1}^{2}\ell_{1}^{4}}{24}+\mathcal{O}(z_{1}^{5}),$ (2.21)
@@
$\ell_{1}(z_{1},a_{0},\theta)$ $=\sqrt{3}-3Q(\theta)z_{1}+\left(\frac{45Q(\theta)^{2}}{2\sqrt{3}}-\frac{9}{8\sqrt{3}}\right)z_{1}^{2}-\frac{\sqrt{3}}{2}a_{0}+\mathcal{O}\left(|z_{1}^{3}|+|a_{0}z_{1}|+|a_{0}^{2}|\right).$
Again, we can use this expression for $\ell_{1}(z_{1},a_{0},\theta)$ to solve for $a_{0}$ using the average-density condition (2.5). Substituting, we obtain
$0$ $=2\sqrt{3}\pi^{2}-\left(\iint_{\Omega}\left(A_{0}+A_{1}V(x,y)\right)dydx-\iint_{\Omega_{0}^{c}}\left(A_{0}+A_{1}V(x,y)\right)dydx\right)$
$=-2\sqrt{3}\pi^{2}\cdot a_{0}+\iint_{\Omega_{0}^{c}}\left(\left(1+a_{0}\right)+\frac{2}{3}(1+a_{0}+z_{1}^{2})\left(2\cos(x)\cos\left(\frac{y}{\sqrt{3}}\right)+\cos\left(\frac{2y}{\sqrt{3}}\right)\right)\right)dydx$
$=-2\sqrt{3}\pi^{2}\cdot a_{0}+2\int_{0}^{2\pi}\int_{0}^{z_{1}\ell_{1}(z_{1},a_{0},\theta)}\left[(1+a_{0})\frac{r^{2}}{3}+z_{1}^{2}\left(-1+\frac{r^{2}}{3}\right)+A_{1}Q(\theta)r^{3}+\mathcal{O}(r^{4})\right]r\ drd\theta$
$=-2\sqrt{3}\pi^{2}\cdot a_{0}-3\pi(1-a_{0})z_{1}^{4}+\mathcal{O}\left(z_{1}^{6}\right),$
@@
Lastly, we use (2.4) to relate these quantities to $\widetilde{\mu}$}:
$0=$ $V(x,y)A_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega-\Omega_{0}^{c}}V(x-\xi,y-\eta)\left(A_{0}+A_{1}\cdot V(\xi,\eta)\right)d\xi d\eta$
$=$ $V(x,y)\left(A_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(A_{1}\sqrt{3}\pi^{2})\right)$
$+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}V(x-\xi,y-\eta)\left(A_{0}+A_{1}\cdot V(\xi,\eta)\right)d\xi d\eta.$
Equating coefficients of $V(x,y)$, using polar coordinates for the integral over the bubbles, we find
$0=$ $-\sqrt{3}{\widetilde{\mu}}\pi^{2}A_{1}+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}\frac{1}{3}V(\xi,\eta)(A_{0}+A_{1}V(\xi,\eta))d\xi d\eta$
$=$ $-\sqrt{3}{\widetilde{\mu}}\pi^{2}A_{1}+2\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\int_{0}^{2\pi}\int_{0}^{z_{1}\ell_{1}}\left(\frac{1}{2}z_{1}^{2}-\frac{1}{6}r^{2}-\frac{Q(\theta)}{3}r^{3}+\mathcal{O}(r^{4},r^{2}z_{1}^{2})\right)r\ drd\theta$
@@
As in the square symmetry case for bubbles, we find that $z_{1}$ scales like ${\widetilde{\mu}}^{{1}/{4}}$, and we obtain a unique $z_{1}=z_{1}({\widetilde{\mu}}^{{1}/{4}})$, for ${\widetilde{\mu}}^{1/4}$ sufficiently small, by the implicit function theorem, with
$z_{1}=\left(\frac{4\pi^{3}}{3}{\widetilde{\mu}}\right)^{1/4}+\mathcal{O}({\widetilde{\mu}}^{1/2}).$ (2.23)
Using the relation $\ell=\sqrt{3}z_{1}+\mathcal{O}(z_{1}^{2})$, this gives us the expression in Theorem 2 as desired.
@@ 2.3.2 Proof of Theorem 2 (iii) — hexagons
We now turn to the case $A_{1}<0$. and proceed analogously. We scale
$A_{0}$ $=1+a_{0},\qquad A_{1}=-\frac{1}{3}\left(1+a_{0}+z_{1}^{2}\right),\qquad a_{0}=a_{0}(z_{1}),$
for a small parameter $z_{1}$. Setting the solution equal to 0 in the free-boundary equation (2.6) and expanding in polar coordinates yields
@@
$0$ $=2\sqrt{3}\pi^{2}-\iint_{\Omega}\left(A_{0}+A_{1}V(x,y)\right)dydx+\iint_{\Omega_{0}^{c}}\left(A_{0}+A_{1}V(x,y)\right)dydx$
$=-2\sqrt{3}\pi^{2}\cdot a_{0}+\iint_{\Omega_{0}^{c}}\left(\left(1+a_{0}\right)-\frac{1}{3}(1+a_{0}+z_{1}^{2})\left(2\cos(x)\cos\left(\frac{y}{\sqrt{3}}\right)+\cos\left(\frac{2y}{\sqrt{3}}\right)\right)\right)dydx$
@@
$0=$ $V(x,y)A_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega-\Omega_{0}^{c}}V(x-\xi,y-\eta)\left(A_{0}+A_{1}\cdot V(\xi,\eta)\right)d\xi d\eta$
$=$ $V(x,y)\left(A_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(A_{1}\sqrt{3}\pi^{2})\right)$
$+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}V(x-\xi,y-\eta)\left(A_{0}-A_{1}V(\xi,\eta)\right)dxdy.$
Equating coefficients of $V(x,y)$, using polar coordinates for the integral over the bubble, we calculate
$0=$ $-\sqrt{3}{\widetilde{\mu}}\pi^{2}A_{1}+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}^{c}}\frac{1}{3}V(\xi,\eta)(A_{0}+A_{1}V(\xi,\eta))d\xi d\eta$
@@
$=$ $\frac{{\widetilde{\mu}}\pi^{2}}{\sqrt{3}}\left(1+z_{1}^{2}-\frac{\sqrt{3}}{4\pi}z_{1}^{4}\right)+\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\left(-\frac{3\pi}{2}z_{1}^{4}+\mathcal{O}(z_{1}^{6})\right).$
Again, we find that $z_{1}$ scales like ${\widetilde{\mu}}^{1/4}$, and we obtain a unique $z_{1}=z_{1}({\widetilde{\mu}}^{1/4})$, for ${\widetilde{\mu}}^{1/4}$ sufficiently small, by the implicit function theorem, with
$z_{1}=\left(\frac{2\pi^{3}}{3}{\widetilde{\mu}}\right)^{1/4}+\mathcal{O}({\widetilde{\mu}}^{3/4}).$ (2.28)